# Autonomous Behavior and Whole-Brain Dynamics Emerge in Embodied Zebrafish Agents with Model-based Intrinsic Motivation

### v1

@@ 3 Methods
In classical reinforcement learning, policies are optimized by mapping a desired behavior to maximizing a reward function [48]. This function is one component of the Markov Decision Process (MDP) that formally specifies a task $M$ as a tuple $\left(\mathcal{S},\mathcal{A},T,p_{0},r\right)$, where $\mathcal{S}$ is the space of environment states, $\mathcal{A}$ is the space of actions, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{S})$ is the transition dynamics (where $\mathcal{P}(\mathcal{S})$ is the set of probability densities over $\mathcal{S}$), $p_{0}$ is the distribution over initial states, and $r$ is the reward function. Importantly, when reward is defined as part of the task MDP, it is extrinsic and provided from the environment. However, learning complex behaviors using extrinsic reward can fail for many real-world behaviors where $r$ is extremely sparse (winning a long-horizon game like Go) or intractable to begin with (intellectual pursuits such as knowledge acquisition) without a powerful exploration mechanism that guides behavior. Model-based intrinsic motivation is a class of such mechanisms that leverage internal world models to convey exploration-relevant information in the form intrinsic reward $r_{t}^{i}$. In the absence of any extrinsic reward $r_{t}^{e}$, policy learning with intrinsic motivation is completely self-supervised. In this work, we consider world models $\omega$ that take the form of forward dynamics prediction. Let $\theta$ parameterize a neural network and $\omega_{\theta}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(S)$. For simplicity, we assume a fixed variance Gaussian density $\omega_{\theta}=\mathcal{N}\left(\phi(\mathbf{s}_{t+1})\mid\phi(\mathbf{s}_{t}),\mathbf{a}_{t};\mu_{\theta},\sigma I\right)$.
Our Intrinsic Drive Module (2A) implements four alternative model-based intrinsic motivation methods as alternative objectives. The Intrinsic Curiosity Module (ICM) [3] defines intrinsic reward as the Shannon surprise of the forward model, $r_{t}^{i}:=-\log\omega_{\theta}(\phi^{I}_{t}\mid\phi_{t},\mathbf{a}_{t})$. $I$ denotes an augmented inverse feature space that is learned on-top of sensory embeddings from $\phi$—ICM trains an additional embedding layer $\phi^{I}$ using an inverse dynamics model parameterized by $\theta_{I}$ by optimizing an MSE loss $L(\theta_{I})=\mathbb{E}_{\pi_{\theta}}\left\|f(\theta_{I};\phi^{I}_{t},\phi^{I}_{t+1})-\mathbf{a}_{t}\right\|_{2}^{2}$. Random Network Distillation (RND) [4] defines a fixed random nonlinear projection of sensory features $g(\phi)$ and trains a predictor network $\hat{g}$ using an MSE loss $r_{t}^{i}:=L(\theta_{\mathrm{RND}})=\mathbb{E}_{\pi_{\theta}}\left\|g(\phi_{t})-\hat{g}(\theta_{\mathrm{RND}};\phi_{t})\right\|_{2}^{2}$. With the distillation objective as the intrinsic reward, RND does not reinforce behaviors by scoring their predictability by a forward dynamics model as in ICM; instead, the random memory provides a simple exploration bonus for visiting novel states under the policy distribution. Disagreement [5] learn an ensemble of world models $\{\omega_{\theta_{j}}\}_{j=1}^{N}$ and defines intrinsic reward as $r_{i}^{t}:=\mathrm{Var}\left(\{\omega_{\theta_{j}}\}_{j=1}^{N}\right)$ for $N$ randomly initialized world models. Ensemble variances scores the stochasticity of the environment and reinforces state-action pairs for which models disagree. Finally, $\gamma$-Progress [7] leverages the temporal history of Shannon surprise to define intrinsic reward using prediction gain, $r_{i}^{t}:=\log\frac{\omega_{\theta_{new}}}{{\omega_{\theta_{old}}}}$, where $\theta_{new}$ parameterized a world model after learning on new transitions withheld from an lagging model $\theta_{old}$.
@@
Autonomous animal behavior enables intelligent and robust decision-making in complex environments. Thus, an intrinsic exploration mechanism that captures the ethology of animal autonomy should satisfy at least two desiderata. First, animals do not perseverate on random noise or pursue stimuli they cannot causally interact with, so animal autonomy should avoid unpredictable or uncontrollable stimuli. For example, Mu et al. [2] demonstrate that zebrafish transition to passive behavioral states when motor commands elicit unpredictable sensory-feedback (unlearnable) or when sensory-feedback is withheld altogether (uncontrollable). Second, animals exhibit consistent rational decision-making strategies across repeated encounters of the same context, so animal autonomy should converge to reliable and stable behavioral responses in such situations. For example, zebrafish exhibit stable behavioral state-switching in the open-loop experimental protocol across trials and subjects [2]. Existing intrinsic drives that rely on prediction-error alone (ICM, RND) perseverate on unlearnable and uncontrollable stimuli by reinforcing stochastic transition dynamics. Learning progress and disagreement overcome this by leveraging temporal dynamics or a world-model ensemble, but together with ICM and RND, suffer from non-stationary rewards—repeated behavioral strategies cause intrinsic reward to saturate such that the reinforcement signal for any single strategy is transient. Although these properties are suitable for exploration in many robotics and machine learning problem settings [36, 37, 38, 39], they may fail to capture the nature of autonomous exploration in animals.
In order to overcome these drawbacks, we introduce Model-Memory-Mismatch Progress (3M-Progress), a novel intrinsic drive that incorporates simple normative qualities of animal autonomy inspired by zebrafish behavior. First, we consider perseveration on unpredictable or uncontrollable stimuli. Learning progress leverages temporal dynamics by comparing predictions from an online memory ($\theta_{new}$) and a long-term memory ($\theta_{old}$), where the memories are mathematically coupled by exponential weight decay. We argue that this long-term memory need not be implemented by decaying the online memory itself—instead, we can decouple these memories completely by learning $\theta_{old}$ in a different context and freezing it during learning $\theta_{new}$. Then, the predictions between memories encode a structural bias towards the environment dynamics from the old context, which can then be leveraged to partition the state-action space in the new context into regions where memories systematically agree or systematically disagree.
Formally, let $\omega_{\theta_{{old}}}(\bar{s}_{t+1}\mid\bar{s}_{t},a_{t})$ parameterize a transition density over state-action space $\bar{\mathcal{S}}\times\mathcal{A}$ from an prior MDP context ¯$\bar{\mathcal{M}}=\left(\bar{\mathcal{S}},{\mathcal{A}},\bar{T},\bar{p}_{0},\bar{r}\right)$. As $L_{\theta_{old}}$ is minimized, the predicted next state $\mathbb{E}_{\omega_{\theta_{{old}}}}\left[\bar{\mathbf{s}}_{t+1}\mid\bar{\mathbf{s}}_{t},\mathbf{a}\right]$ approaches the true transition $\bar{\mathbf{s}}_{t+1}\sim\bar{T}\left[\bar{\mathbf{s}}_{t+1}\mid\bar{\mathbf{s}}_{t},\mathbf{a}_{t}\right]$. In the new MDP context $\mathcal{M}$, the same property holds, where $\mathbb{E}_{\omega_{\theta_{{new}}}}\left[{\mathbf{s}}_{t+1}\mid{\mathbf{s}}_{t},\mathbf{a}_{t}\right]\rightarrow{\mathbf{s}}_{t+1}$ as $L_{\theta_{\mathrm{}new}}\rightarrow 0$. Then, if there exists a state-action pairs $(\bar{\mathbf{s}}_{t},a_{t})\in\bar{\mathcal{S}}\times\mathcal{A}$ and $({\mathbf{s}}_{t},a_{t})\in{\mathcal{S}}\times\mathcal{A}$ where $\mathcal{D}_{KL}\left[\bar{T}(\bar{\mathbf{s}}_{t+1}\mid\bar{\mathbf{s}}_{t},\mathbf{a}_{t})\mid\mid T({\mathbf{s}}_{t+1}\mid{\mathbf{s}}_{t},\mathbf{a}_{t})\right]\approx 0$, then
@@
$\epsilon_{t}:=\mathcal{D}_{KL}\left[\omega_{\theta_{{old}}}(\bar{\mathbf{s}}_{t+1}\mid\bar{\mathbf{s}}_{t},\mathbf{a}_{t})\mid\mid\omega_{\theta_{{new}}}(\bar{\mathbf{s}}_{t+1}\mid\bar{\mathbf{s}}_{t},\mathbf{a}_{t})\right]\propto\left\|\mu_{\theta_{{new}}}-\mu_{\theta_{{old}}}\right\|_{2}^{2}\approx 0.$
In other words, state-action pairs with similar environment dynamics result in similar world-model predictions. We term $\epsilon_{t}$ model-memory-mismatch, which loosely partitions the state-action space into model-memory agreement ($\epsilon_{t}=0$) and disagreement ($\epsilon_{t}\gg 0$). Depending on the transition dynamics in each context, model-memory-mismatch encodes distinct behavioral states.
Figure 2B illustrates a specific example inspired by zebrafish, where $\theta_{old}$ is learned in an ethological environment—the agent can freely behave and experiences passive fluid forces as it swims and learns to positionally homeostate by resisting an opposing current induced by visual flow. Importantly, the reward in this context is shaped such that the agents rests once it regains its position in order for $\omega_{\theta_{{old}}}$ to learn a suitable distribution of state-action transitions from both active and passive behaviors. Then, the agent is put into an unethological experimental protocol that headfixes the agent and swim commmands do not elicit sensory feedback. As the agent behaves in this environment, it constructs a new memory $\omega_{\theta_{{new}}}$. The intrinsic reward take a similar mechanistic form as learning progress but filters model-memory-mismatch instead of memory parameters,
$r_{t}^{i}=f(\hat{\epsilon}_{t}-\epsilon_{t}),$
where $\hat{\epsilon}_{t}=(1-\gamma)\hat{\epsilon}_{t-1}+\gamma\epsilon_{t}$ and $f$ is an activation function. Thus, the exponential filter avoids perseveration on behaviors where memories disagree when $f$ is an identity operator. In order to encourage periodic exploration to behavioral partitions where memories disagree, we can simply take $f$ to be any symmetric function. This reinforces deviations from the moving baseline in either direction. Unlike learning progress, this formulation does not saturate as learning unfolds since $\epsilon_{t}$ is computed using a fixed memory. In fact, stable predictions from $\omega_{\theta_{new}}$ due more environment interactions only increases the signal-to-noise ratio in $\epsilon_{t}$, resulting in more robust behavioral patterns. The shape of $f$ and the progress-horizon $\gamma$ determine the relative time in each partition and periodicity of switching. In all our experiments, we set $f(x)=|x|$ and $\gamma=0.99$. Additionally, all experiments include an action penalty $r^{a}_{t}=-\lambda\|\mathbf{a}_{t}\|_{2}^{2}$ to encourage exploration of passive behavior during policy learning.
Owing to the formulation of 3M-Progress, learning the state-value function requires, in principle, the presence of units in the core module that are functionally equivalent to neurons in the Noradernergic cluster of the Medulla Oblongata (NE-MO) and radial astrocytes identified in zebrafish by Mu et al. [2] (2A). 3M-Progress detects sensory-motor mismatch using a prior memory as an expectation of how action is coupled to sensory-feedback under normal, ethological environment dynamics (2C). This is functionally equivalent to signaling failed swim-attempts by NE-MO neurons in zebrafish. Such a signal directly implicates a biological prior that encodes how swim-attempts should be coupled to visual feedback. Similarly, the exponential weighted average is a discrete-time leaky integrator on model-memory-mismatch (NE-MO input), which is functionally equivalent to radial astrocytes that accumulate NE-MO signals during failed swim-attempts and decay during passivity (2C).
3.1 Model-Brain and Inter-Animal Alignment
To rigorously evaluate our embodied agents against whole-brain response patterns collected by [2], neural-glial recordings from zebrafish were first aligned to periods specifically capturing behavioral state transitions between active swimming and futility-induced passivity. Correspondingly, latent states and predicted neural-glial responses from our agents were aligned to these same experimentally defined epochs using event-triggered averaging around transition events. To quantify model-brain alignment, we employed a stringent “One-to-One” mapping, in which each neural or astrocytic unit from the biological data was matched directly to the single most correlated artificial unit from the computational model. Although this mapping is typically too restrictive for capturing inter-animal alignment in heterogeneous neural populations for brain-region-specific (e.g. not whole brain) data in other animals (as shown in prior sensory [49] and cognitive systems [50] where full linear regression is necessary), we found it was fully sufficient here, due to the exceptional whole-brain reliability of zebrafish neural-glial responses during futility-induced passivity. Indeed, inter-animal alignment computed from pairwise correlations between individual zebrafish was nearly 100%, establishing a robust empirical ceiling for evaluating model performance.
Specifically, inter-animal alignment was computed as the correlation of neural-glial response patterns across pairs of zebrafish within aligned epochs, thus formally defining the upper bound of predictivity achievable by any candidate model. This approach ensures that successful computational models achieve neural predictivity indistinguishable from biological measurements, within the bounds set by natural biological variability. Together, these procedures allowed us to precisely quantify model-brain alignment, and achieving close correspondence between artificial and biological responses under this strict One-to-One mapping criterion provides strong evidence that our intrinsic motivation-driven embodied agents effectively capture the detailed neural and glial dynamics underlying autonomous behavioral transitions in zebrafish.
3.2 Model-Behavioral Alignment
To quantify similarity between biological zebrafish and our artificial agents, we use Pearson’s $r$ as a metric between the respective swim power readouts for each system surrounding state transitions. In zebrafish, swim power was calculated as the standard deviation over a 10ms window of recorded tail motor nerve signals following Mu et al. [2]. We use the same passive and active transitions windows identified by Mu et al. [2] to compute the inter-animal behavioral alignment across 11 subjects by applying the metric to swim power after smoothing and normalization over a 20 second window surrounding the transition time. For the artificial agents, we take the norm of their joint torques as swim power and identify behavioral transitions as high-frequency changes in swim power above a threshold of 1 Newton. This threshold was determined empirically by recording joint activations during active and passive behaviors in the default swim task in the dm-control suite [1]. Model-behavioral consistency is then computed by applying the metric between segmented zebrafish data and the agent’s swim power surrounding these transitions over a 20-step window.
4 Results
@@
Figure 4: Model-Brain Alignment. A) Noise-corrected Pearson’s $r$ correlation between whole-brain neural and glial units and artificial units from trained agents during passive transitions. B) Model scores on behavioral and whole-brain alignment during passive transitions.

### v2

@@ 3 Methods
In classical reinforcement learning, policies are obtained by mapping a desired behavior to maximizing a reward function [56]. This function is one component of the Markov Decision Process (MDP) that formally specifies a task $M$ as a tuple $\left(\mathcal{S},\mathcal{A},T,p_{0},r\right)$, where $\mathcal{S}$ is the space of environment states, $\mathcal{A}$ is the space of actions, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{S})$ is the transition dynamics (where $\mathcal{P}(\mathcal{S})$ is the set of probability densities over $\mathcal{S}$), $p_{0}$ is the distribution over initial states, and $r$ is the reward function. Importantly, when reward is defined as part of the task MDP, it is extrinsic and provided from the environment. However, learning complex behaviors using extrinsic reward can fail for many real-world situations where $r$ is extremely sparse (winning a long-horizon game like Go) or intractable to begin with (intellectual pursuits such as knowledge acquisition) without a powerful exploration mechanism to guide behavior. Model-based intrinsic motivation is a class of such mechanisms that leverage predictive world models to convey exploration-relevant information using prediction-error to the form an intrinsic reward $r_{t}^{i}$. In the absence of any extrinsic reward $r_{t}^{e}$, policy learning with intrinsic motivation is completely self-supervised. In this work, we consider intrinsic motivation driven by forward dynamics world models $\omega$. Let $\theta$ parameterize a neural network and $\omega_{\theta}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(S)$. For simplicity, we assume a fixed variance Gaussian density $\omega_{\theta}=\mathcal{N}\left(\phi(\mathbf{s}_{t+1})\mid\phi(\mathbf{s}_{t}),\mathbf{a}_{t};\mu_{\theta},\sigma I\right)$. Here, $\phi(\cdot)$ denotes the concatenated sensory embeddings from $\phi_{I}(\cdot)$ and $\phi_{J}(\cdot)$. This class of methods is appealing from two complimentary perspectives. On one hand, it is computationally straightforward; learning an internal model of an MDP’s transition dynamics is a natural way to measure novelty of the states visited under the agent’s policy—states with a high prediction error under the internal model indicate regions of the state-action space that are poorly understood or rarely visited. On the other hand, it is well motivated by experimental evidence in both neuroscience and psychology; numerous empirical studies suggest humans and animals depend on predictive models of the world for decision-making [57, 58, 59, 60, 61]. Thus, world models are well-positioned as a primary substrate for intrinsic motivation; however, we emphasize that a world model is best viewed as only a substrate, as exactly how it should be used for intrinsic motivation remains unclear.
Model-Brain and Inter-Animal Alignment
To rigorously evaluate our embodied agents against whole-brain data collected by Mu et al. [2], neural-glial recordings from zebrafish were first aligned to periods specifically capturing behavioral state transitions between active swimming and futility-induced passivity. Correspondingly, latent states and predicted neural-glial responses from our agents were aligned to these same experimentally defined epochs using event-triggered averaging around transition events. To quantify model-brain alignment, we employed a stringent “One-to-One” mapping, in which each neural or astrocytic unit from the zebrafish brain was matched directly to the single most correlated artificial unit from the virtual embodied agent. Although this mapping is typically too restrictive for capturing inter-animal alignment in heterogeneous neural populations for brain-region-specific (e.g. not whole brain) data in other animals (as shown in prior sensory [62] and cognitive systems [63] where full linear regression is necessary), we found it was fully sufficient here due to the exceptional whole-brain reliability of zebrafish neural-glial responses during futility-induced passivity. Indeed, inter-animal alignment computed from pairwise correlations between individual zebrafish was nearly 100%, establishing a robust empirical ceiling for evaluating model performance.
Inter-animal alignment was computed as the correlation of neural-glial response patterns across pairs of zebrafish within aligned epochs, thus formally defining the upper bound of predictivity achievable by any candidate model (the mathematical details of the chosen metric can be found in Appendix F). This approach ensures that successful computational models achieve neural predictivity indistinguishable from biological measurements, within the bounds set by natural biological variability. Together, these procedures allowed us to precisely quantify model-brain alignment; achieving close correspondence between artificial and biological responses under this strict One-to-One mapping criterion provides strong evidence that our intrinsic motivation-driven embodied agents effectively capture the detailed neural-glial dynamics underlying autonomous behavioral transitions in zebrafish.
Model-Behavioral Alignment
To quantify similarity between biological zebrafish and artificial agents, we use Pearson’s $r$ as a metric between the respective swim power readouts for each system surrounding state transitions. In zebrafish, swim power was calculated as the standard deviation over a 10ms window of recorded tail motor nerve signals following Mu et al. [2]. We use the same passive and active transitions windows identified by Mu et al. [2] to compute the inter-animal behavioral alignment across 11 subjects by applying the metric to swim power after smoothing and normalization over a 20 second window surrounding the transition time. For the artificial agents, we take the norm of their joint torques as swim power and identify behavioral transitions as high-frequency changes in swim power above a threshold of 1 Newton. This threshold was determined empirically by recording joint activations during active and passive behaviors in the default swim task in the dm-control suite 6-link swimmer environment [1]. Model-behavioral consistency is then computed by applying the metric between segmented zebrafish data and the agent’s swim power surrounding these transitions over a 20-step window.
4 Animal-like Exploration from First Principles: Lessons from Zebrafish
Autonomy in animals enables intelligent and robust decision-making in complex environments, even in the presence of high-entropy or unfamiliar dynamics such as noisy stimuli or laboratory habitats. Reflecting on characteristics of autonomous exploration in zebrafish, we propose two simple desiderata for intrinsic drives that capture animal-like autonomy:
  1. 1.
    Animals do not perseverate on unpredictable stimuli or pursue stimuli they cannot causally interact with. Mu et al. [2] demonstrate that zebrafish transition to passive behavioral states when motor commands elicit unpredictable sensory-feedback (unlearnable dynamics) or when sensory-feedback is withheld altogether (uncontrollable dynamics). An intrinsic objective for animal autonomy should avoid unpredictable or uncontrollable stimuli.
  2. 2.
    Animals exhibit consistent decision-making strategies across repeated encounters of the same context. Zebrafish exhibit stable behavioral state-switching in the open-loop experimental protocol across trials and subjects, ultimately converging on a single exploration policy [2]. An intrinsic objective for animal autonomy should converge to a stable behavioral policy.
Intrinsic objectives that rely on prediction-error alone [3, 4] reward stochastic environment dynamics and incentivize learning transitions in which the agent has no causal control [^2] Although [3] proposed an inverse dynamics feature space to avoid representing uncontrollable stimuli, it’s efficacy has not been demonstrated outside of simple pixel-based environments with discrete actions.. Learning progress [6, 7, 64] and Disagreement [5] overcome this by leveraging temporal dynamics or statistics of a world model ensemble, but together with ICM and RND, are formulated as functions of the world model training loss and are thus non-stationary—learning on repeated behavioral strategies drive the training loss towards zero, and so any single behavioral strategy is transient since it is not consistently reinforced. Although these properties are suitable for exploration in some robotics domains, particularly when the policy is supplemented with an extrinsic task reward [36, 37, 38, 3, 5], they fail to capture the nature of autonomous exploration in animals.
@@ 4 Animal-like Exploration from First Principles: Lessons from Zebrafish
In order to overcome these drawbacks, we introduce Model-Memory-Mismatch Progress (3M-Progress), a novel intrinsic drive that incorporates these simple normative properties of animal autonomy inspired by zebrafish. Curiosity, disagreement, and learning progress couple intrinsic reward to a moving world model, yielding a non-stationary policy objective. The RL problem becomes a two player minimax game, where the actor seeks states that increase prediction-error and the learner reduces it, leading to a reward landscape that flattens and precludes a stationary optimal policy [^3] See Appendix C.1 for mathematical details.. Thus, each of the existing curiosity-driven exploration algorithms we consider fails to converge to an exploratory policy. However, learning progress [6] is unique among these algorithms in its ability to avoid perseveration on unpredictable and uncontrollable stimuli by leveraging the temporal dynamics of prediction-error. This is achieved by maintaining two world models, an online model and a long-term memory implemented as an exponential weight decay on the online model. The slow timescale memory provides a moving baseline that, when compared with the online model, creates an intrinsic reward that flattens as the temporal dynamics of the online model’s predictions flatten—in other words, a computationally efficient and biologically plausible implementation of a time-derivative.
To combine the adaptive behavior afforded by derivative-like operations with an intrinsic goal that admits a stationary solution, we propose to decouple the two memories completely. We achieve this by learning the long-term memory $\omega_{\theta}$ in a pretraining environment and deploying it as a fixed prior while the online model $\omega_{\theta^{\prime}}$ is learned in a new environment with new transition dynamics. The frozen “ethological memory” functions as a static memory primitive: comparing its predictions with the online memory produces a residual that encodes an explicit bias toward transition dynamics that match the pretraining environment, which can then be leveraged to partition the state-action space in the new context into regions where memories systematically agree or disagree  [^4] See Appendix C.2 for mathematical details.. This reflects the idea that animals develop in an ecological niche with characteristic environment dynamics, distilled via experience as an internal world model $\omega_{\theta}$. In new environments with different transition dynamics, such as the experimental protocol in Mu et al. [2], animals can use this prior to seek or avoid regions that match their ecological niche as they learn new world model $\omega_{\theta^{\prime}}$. To capture this, we define the model-memory-mismatch ($\epsilon_{t}$), exponential filter ($\hat{\epsilon}_{t}$), and niche-aware intrinsic motivation $r_{t}^{i}$ as
@@
$\epsilon_{t}$ $:=\mathcal{D}_{\mathrm{KL}}\left[\omega_{\theta}\left(\phi({\mathbf{s}}_{t+1})\mid\phi({\mathbf{s}}_{t}),\mathbf{a}_{t}\right)\mid\mid\omega_{\theta^{\prime}}\left(\phi({\mathbf{s}}_{t+1})\mid\phi({\mathbf{s}}_{t}),\mathbf{a}_{t}\right)\right],$ (1)
$\hat{\epsilon}_{t}$ $=(1-\gamma)\hat{\epsilon}_{t-1}+\gamma\epsilon_{t},$ (2)
$r_{t}^{i}$ $=|\hat{\epsilon}_{t}-\epsilon_{t}|,$ (3)
where $\mathcal{D}_{\mathrm{KL}}$ denotes the Kullback-Leibler divergence and $\gamma$ is the filter timescale. The non-negativity of this divergence coarsely partitions the state-action space into niche-seeking (model-memory agreement: $\epsilon_{t}\approx 0$) and niche-avoidance (model-memory disagreement: $\epsilon_{t}\gg 0$). We design the intrinsic reward mechanistically after $\gamma$-Progress [6] to maintain a moving baseline, but filter model-memory-mismatch rather than model parameters.
@@
Figure 4: Model-brain alignment averaged across active and passive transitions. A) Noise-corrected Pearson’s $r$ correlation between whole-brain neural and glial units and artificial units from trained agents. B) Model scores on behavioral and whole-brain alignment.
Due to the moving baseline, the reward does not perseverate on any single partition, either agreement or disagreement. The symmetry enforced by the absolute value encourages periodic exploration between partitions by reinforcing deviations from the moving baseline in either direction. Unlike learning progress, this formulation does not saturate as learning unfolds since $\epsilon_{t}$ is computed using a fixed memory. In fact, as the prediction error from $\omega_{\theta^{\prime}}$ stabilizes with more environment interactions, the signal-to-noise ratio in $\epsilon_{t}$ only increases, resulting in more robust behavioral patterns. Note that the absolute value is a specific choice of an activation function. The shape of this function and the progress horizon $\gamma$ determine the relative time spent in each partition. All experiments and baseline algorithms include an action penalty $r^{a}_{t}=-\lambda\|\mathbf{a}_{t}\|_{2}^{2}$ to encourage exploration of passive behavior. Figure 2B illustrates a specific example inspired by zebrafish, where $\omega_{\theta}$ is learned in an ethological environment—the agent can freely behave and experiences passive fluid forces induced by self-motion as it swims and and active fluid forces as it learns positional homeostasis by resisting an opposing current. The agent is then put into an unethological experimental protocol where it is head-fixed and its swim commands do not elicit sensory feedback. As the agent behaves in this environment, it distills a new memory $\omega_{\theta^{\prime}}$ from experience.
3M-Progress as a Normative Model of Neural-Glial Computation
Owing to the formulation of 3M-Progress, the state-value function must implement units in the core module that are functionally equivalent to neurons in the Noradernergic cluster of the Medulla Oblongata (NE-MO) and radial astrocytes identified by Mu et al. [2] (2A). 3M-Progress detects sensory-motor mismatch using a prior memory as an expectation of how action is coupled to sensory-feedback under ethological environment dynamics (2B). This is functionally equivalent to signaling failed swim-attempts by NE-MO neurons in zebrafish. Similarly, the exponential filter is a discrete-time leaky integrator on model-memory-mismatch (NE-MO input), which is functionally equivalent to radial astrocytes that accumulate NE-MO signals during failed swim-attempts and decay during passivity (2C).
5 Experimental Results